OpenStudy (stamp):

[CALCULUS III - DOUBLE INTEGRALS] Evaluate \[\int_{0}^{1}\int_{y-1}^{0}e^{x+y}dxdy\].

OpenStudy (stamp):

\[\int_{y-1}^{0}e^{x+y}dx=e^{x+y}|_{y-1}^{0}=e^{(0)+y}-(e^{(y-1)+y})\]\[=e^y-e^{2y-1}\]

OpenStudy (stamp):

\[=e^y(1-e^{y-1})\]

OpenStudy (stamp):

\[\int_{0}^{1}e^y-e^{2y-1}dy\]

OpenStudy (stamp):

Currently I am approaching it by solving the first integral, solving the second integral, then evaluating the addition. But I feel like there is some property I am missing that could allow me to avoid all of that, notice how the limits change.

OpenStudy (stamp):

I did a wolfram computation and have verified that they do cancel out, the first integral is c and the second integral is -c, so you obtain c + -c = 0 Now rather than evaluate each integral, what property or observation am I missing that will allow me to realize the symmetry that allows them to cancel out

OpenStudy (stamp):

properties of double integrals: http://tutorial.math.lamar.edu/Classes/CalcIII/DIGeneralRegion.aspx

OpenStudy (stamp):

Anybody capable and willing to help provide analysis as to how these integrals evaluated and added together equate to zero?

OpenStudy (stamp):

limits flipped gives\[e^{2y-1}-e^y\]

OpenStudy (stamp):

You can factor out the negative to obtain the same integral \[e^y-e^{2y-1}\]

OpenStudy (stamp):

So you are taking \[\int_{0}^{1}u\ du+(-\int_{0}^{1}u\ du)\]\[\int_{0}^{1}u\ du-\int_{0}^{1}u\ du=0\] Thanks all